OpenStudy (anonymous):

A concert hall earned $40,000 in sold−out comedy ticket sales. Front section tickets cost $150 per seat and back section tickets cost $50 per seat. The number of front section seats is three times the number of back section seats. How many seats are in the front section?

OpenStudy (anonymous):

150x+50y=40000 3x=y

OpenStudy (anonymous):

my english is bad may be 3y=x

OpenStudy (anonymous):

yeah it is \(x=3y\)

OpenStudy (anonymous):

\[150x+50y=40000\] \[x=3y\] \[150\times 3y+50y=4000\] \[450y+50y=4000\] \[500y=400\] etc

OpenStudy (anonymous):

Thanks